Which of the following describes a scenario in which the molecular geometry would have the smallest bond angle(s)?

  • four bonding domains (or bonding pairs) and zero nonbonding domains (or lone pairs)

  • two bonding domains (or bonding pairs) and two nonbonding domains (or lone pairs)

  • three bonding domains (or bonding pairs) and one nonbonding domain (or lone pair)

  • two bonding domains (or bonding pairs) and zero nonbonding domains (or lone pairs)

  • two bonding domains (or bonding pairs) and one nonbonding domain (or lone pair)

  • three bonding domains (bonding pairs) and zero nonbonding domains (or lone pairs)
Answer #1

#Bonding pairs=BP, # Lone pair = LP

1) 4-BP and 0-LP = Tetrahedral shape and bond angle =109.5'

2) 2-BP and 2- LP = Bent shape and bond angle = 104.5'

3) 3-BP and 1-LP = Trigonal Pyramidal shape and bond angle =107'

4)2-BP and 0-LP = Linear shape and bond angle =180'

5) 2-BP and 1-LP = Bent shape and bond angle =120'

6)3-BP and -LP = Trigonal planar shape and bond angle =120'

From this clearly can see that 2-BP and 2- LP will be having smallest bond angle.